OpenStudy (anonymous):

What is the sum of a 52-term arithmetic sequence where the first term is 6 and the last term is 363?

OpenStudy (kirbykirby):

So, if you have an arithmetic sequence, it has the form: \(a, a+d, a+2d, a+3d, \ldots, a+nd\) Where \(a\) is the first term, and \(d\) is the common difference. Here, there are \(n+1\) terms in the sequence (not \(n\) terms), since the first term does not have a "\(d\)" in it. Here, \(a=6\). Now, you can write the last term (the 52-nd term) written generally as \(a+nd\), more specifically as \(6+51d\) since you need 52 terms in total. The last term is 363, so \(363=6+51d\implies d=7\). Now, adding up the terms of your sequence: \(6 + (6+7) + (6+2\cdot 7) + \ldots +6+51\cdot 7\) To sum these 52 terms in sigma (summation) notation, you essentially do: \[\large \sum_{d=0}^{52-1}(6+d\cdot 7)\] For this, you need the formula: \[\large \sum_{d=0}^nd=\frac{n(n+1)}{2}\] Let me know if you need more help.